OpenStudy (anonymous):

solve for x 4(2x – 5) = 7 – x

OpenStudy (godlovesme):

Distribute 4 first

OpenStudy (idealist10):

4(2x-5)=7-x 8x-20=7-x -20=7-x-8x -20=7-9x 9x=7-(-20) 9x=7+20 9x=27 x=27/9 x=3 Answer: 3